# Vertex Fitting In Low-Material Budget Pixel Detectors

### v1

@@ Abstract
This paper provides a detailed description of a vertex fitting algorithm suitable for precision measurements in low-energy particle physics experiments. An accurate reconstruction of low-momentum trajectories can be accomplished by reducing the material budget of the detector to a few per mill of the radiation length. This limits the multiple scattering undertaken by particles inside the detector and improves the vertex fitting accuracy. However, for sufficiently light detection systems, additional sources of errors, such as the intrinsic spatial resolution of the sensors, must be considered in the reconstruction of the vertex parameters. The algorithm developed in this work provides a complete treatment of multiple scattering and spatial resolution in the context of vertex fitting for light pixel detectors. In addition to this, a study of the vertex reconstruction in the low-material budget pixel detector of the Mu3e experiment is presented.
I Introduction
@@
With the increase of the instantaneous and integrated beam luminosities, the requirements of particle physics experiments for precise tracking and vertexing detectors, with high radiation tolerance, have become more stringent e.g., SNOEYS2023168678 ; Hartmut2018 ; CARNESECCHI2019608 ; MOSER201685 . In this regard, silicon pixel sensors can provide high granularity, low material budget structures and the radiation-hardness that most experiments need, e.g., MOSER201685 ; SPANNAGEL2019612 ; Abelev_2014 ; ARNDT2021165679 . It is important, however, that precise detection systems are developed in conjunction with equally performing analysis methods for the reconstruction of particle trajectories and decay vertices, e.g. FRUHWIRTH1987444 ; BILLOIR1992139 ; Waltenberger2007 ; RevModPhys.82.1419 . To this aim, several fitting algorithms have been designed and optimized over the years to deal with hit selection, pattern recognition, errors calculations and high track multiplicity (see for instance Mankel_2004 ; RevModPhys.82.1419 and references therein). The practical implementation of these methods must be tailored around the actual detector and magnetic field configuration of the experiments. This makes tracking and vertexing a topic which is in continuous evolution adapting itself to the new challenges set by upcoming experiments.
@@
This study addresses the problem of vertex fitting in the low-material budget pixel detector of the Mu3e experiment ARNDT2021165679 . As explained in section II, the detector design has been optimized to minimize Multiple Coulomb Scattering (MS) effects on particle trajectories and signal kinematics. However, for light detectors such as Mu3e, the intrinsic pixel resolution becomes another limiting factor in the vertex reconstruction which cannot be ignored. Under these circumstances, the vertex fitting should account for MS and pixel resolution in the error calculations as well as for any energy loss in the detector which may hamper tracks momentum reconstruction. The present algorithm deals with all these sources of errors and it is illustrated in section III whilst in section IV, a comparative study is made among different inclusive scenarios that encompass: (A) MS only; (B) MS and pixel resolution together; (C) all sources of errors (MS, pixel resolution and energy losses) are included.
II The low-material budget Mu3e pixel detector
@@
at Branching Ratios (BR) $>10^{-16}$ ARNDT2021165679 . This threshold is $4$ orders of magnitude smaller than previous experimental upper limits ($10^{-12}$) BELLGARDT19881 and $38$ orders of magnitude larger than theoretical Standard Model (SM) calculations (BR $=10^{-54}$), e.g., MARCIANO1977303 ; RevModPhys.73.151 . However, new theoretical models predict the existence of extra degrees of freedom beyond the SM which may bring CLF violation within the reach of near future experiments such as Mu3e, e.g., KAKIZAKI2003210 ; DEGOUVEA201375 . Consequentially, an observation of $\mu^{+}\rightarrow e^{+}e^{-}e^{+}$ at single event sensitivities aimed by the Mu3e experiment will imply scenarios of new physics.
The process in (1) yields a relatively simple decay topology with the 3 final state leptons produced at the same vertex of interaction and momentum vectors, $\vec{p}$, determined by the energy and momentum conservation for decaying muons at rest. The main background processes in Mu3e measurements are the muon internal conversion $\mu^{+}\rightarrow e^{+}e^{-}e^{+}+\nu_{e}+\bar{\nu}_{\mu}$ (BR $\approx 10^{-5}$) and the combination of one electron and 2 positrons from independent sources, e.g., one Bhabha electron plus two Michel positrons $\mu^{+}\rightarrow 2\times(e^{+}+\nu_{e}+\bar{\nu}_{\mu})+e^{-}$ BELLGARDT19881 .
Figure 1: Scheme of the central barrel of the Mu3e pixel detector, side and x-y views ARNDT2021165679 .
The aimed single event sensitivity of $2\cdot 10^{-15}$, during phase I of the experiment, can be achieved with an energy-momentum resolution of $\lesssim 1$ MeV/c and by using precise vertexing and timing systems ARNDT2021165679 .
@@
Figure 1 shows the schematic of the foreseen Mu3e tracker central station which is important for vertex fitting and track reconstruction ARNDT2021165679 . Two recurl stations, one up-stream and one down-stream, will also be part of the final detector design. These increase the angular acceptance of the experiment and allow to measure long-armed trajectories to achieve improved momentum resolution.
The layers have cylindrical symmetry and are concentrically placed around the target, a hollow double cone made of Mylar 100 mm in length and with a base radius of 19 mm. The target is placed in a solenoid magnetic field of 1 T with the base at a minimum distance of $\approx 4$ mm from the innermost layer of the pixel tracker. Particle trajectories bend inwards following helical trajectories around the field lines possibly making multiple re-curls. Each layer of the pixel detector is sectioned in sub-elements called ladders. A ladder is a series of chips mounted on the same flex-tape. There are 8, 10, 24 and 28 ladders for layer 1, 2, 3 and 4, respectively. For instance, the innermost layer of the tracker, crucial for vertex fitting, is made of 8 ladders each one tilted by a ${}^{\circ}$ angle with respect to the neighbours. This configuration forms a 8-sided surface which extends for $\sim 12$ cm or 6 chips length, see figure 1.
@@
III Vertex fitting in the Mu3e detector
Vertex reconstruction can be accomplished in two steps: vertex finding and vertex fitting, see e.g., RevModPhys.82.1419 . The former consists in grouping trajectories that have been most likely produced in the same decay process. The latter involves finding the most likely vertex coordinates $(x,y,z)_{\text{v}}$ and the initial momentum vectors of all clustered tracks. In Mu3e, vertex finding is accomplished by considering all possible combinations of two positive and one negative tracks in the detector within time frames of 64 ns. For the vertex fitting, a least-squares optimization algorithm has been developed based on the method illustrated in BILLOIR1992139 .
III.1 Track parameters and uncertainties
Trajectories are defined by 6 parameters ($[x,y],z,\phi,\lambda,k)$. These are the coordinates of one point along the track, the angles $\phi$ and $\lambda$ defining the direction of the tangent vector to the trajectory and the factor $k=(p/q)^{-1}$ where $q$ is the charge of the particle and $p$ is the magnitude of the momentum vector. The following relationships among the momentum components in the global Cartesian frame and the angles $\phi$ and $\lambda$ hold true:
@@ III.1 Track parameters and uncertainties
The input measurements to fit are the values of the track parameters at a given Reference Surface (RS) along with their covariance matrix ($\Xi$): ($[x,y],z,\phi,\lambda,k,\Xi)_{\text{meas}}$. In this study, the RS is the innermost layer of the pixel detector which is the closest one to the expected real vertex position. The coordinates $x,y$ are not independent and can be given in a single expression by using a local reference frame with center in the middle of the pixel area $(\bar{x},\bar{y},\bar{z})$ and base vectors $(\hat{u},\hat{z^{\prime}})$. The vector $\hat{z^{\prime}}$ is parallel to the global coordinate $\hat{z}$ while $\hat{u}$ is perpendicular to it and parallel to the pixel surface, see figure 2. The equations that link the local coordinates to the global ones are:
@@
where the angle $\gamma$ is the orientation of the pixel with respect to the $x$ axis of the global reference frame. Following the equations in 3, the track parameters at the RS become $(u,z,\phi,\lambda,k)$  [^1] The coordinate $z^{\prime}$ can be replaced by $z$ given that the constant $\bar{z}$ in eq. 3 does not contribute to the calculations of the derivatives and residuals carried out in section III.2.
Figure 2: Sketch of the RS in the Mu3e vertex fitting. The axes of the global (local) reference frame are drawn in black (green); the $\hat{z^{\prime}}$ axis perpendicular to the $x-y$ plane is not shown. The parameters (${x,y},z)_{\text{meas}}$ are placed in the geometrical centre of the pixel surface $\square$. The fit parameters ($u,z,\phi,\lambda,k)_{\text{fit}}$ are obtained by forward propagating vertex parameters ($x,y,z,\phi,\lambda,k)_{\text{v}}$ via a function $h=h(\textbf{v}=(x,y,z)_{\text{v}},\textbf{t}(\phi_{\text{v}},\lambda_{\text{v}}),k_{\text{v}}))$.
The covariance matrix can be written as the sum of two terms, i.e., $\Xi_{\text{meas}}=\Xi_{\text{track}}+\Xi_{\text{phys}}$. The term $\Xi_{\text{track}}$ accounts for the uncertainties accumulated during track fitting while $\Xi_{\text{phys}}$ accounts for MS and pixel resolution at the RS. Pixel resolution contributes to the smearing of the hit position by $\sigma_{l}=l/\sqrt{12}$, where $l$ is the length of the pixel side. The MS changes the direction of the track upon crossing the RS. The tilt can be approximated by a Gaussian distribution with zero mean and standard deviation $\theta_{\text{MS}}$ Yao_2006 :
$\theta_{\text{MS}}=13.6\,\left[\text{MeV/c}\right]\,\frac{q}{p}\sqrt{\frac{x}{\beta^{2}\text{X}_{o}}}\left(1+0.038\,\text{ln}\left(\frac{d}{\text{X}_{o}}\right)\right)\,.$ (4)
In the previous equation $\beta$ is the relativistic velocity and $d$ is the distance travelled by the particle in the RS. From eq. 4, the changes in $\phi$ and $\lambda$ due to MS, $\sigma_{\phi}$ and $\sigma_{\lambda}$, can be obtained by projecting $\theta_{\text{MS}}$ onto the transverse and longitudinal planes, respectively, see e.g., VALENTAN2009728 :
@@
In addition to MS and pixel resolution, an error $\sigma_{k}$ on the track curvature (or $k$) is introduced in $\Xi_{\text{meas}}$ to account for the energy lost by electrons and positrons in the detector, see e.g., Mankel_2004 . Although energy losses in a thin silicon layer like the RS are negligible RevModPhys.60.663 , they are nevertheless included in the present vertex fitting algorithm to provide a complete treatment of the errors. Without loss of generality, the covariance matrix in this study is written as $\Xi_{\text{meas}}$ = diag$\left\{\sigma_{u}^{2},\sigma_{z}^{2},\sigma_{\phi}^{2},\sigma_{\lambda}^{2},\sigma_{k}^{2}\right\}$ with $\sigma_{u}=\sigma_{z}=\sigma_{l}$.
III.2 Least-squares algorithm
In the vertex fitting, a map that links the track vertex parameters ($x,y,z,\phi^{i},\lambda^{i},k^{i})_{\text{v}}$ to the parameters at the RS ($u^{i},z^{i},\phi^{i},\lambda^{i},k^{i})_{\text{fit}}$ is needed, as shown in figure 2. In what follows, $i=1,2,3$ for the ($e^{-},e^{+},e^{+}$) in a Mu3e decay. Starting from the analytical expression of a helical trajectory of a charged particle in a magnetic field, this map can be written as: $h_{j}=h(\textbf{v},\textbf{t}^{i},k^{i}_{\text{v}})_{j}$, where $h_{j}=(u,z,\phi,\lambda,k)_{\text{fit}}$ for $j=1,2,...,5$, $\textbf{v}=(x,y,z)_{\text{v}}$ and $\textbf{t}=(\phi,\lambda)_{\text{v}}$, see section VII for the analytic expressions of $h_{j}$. In first approximation, this function can be linearized near some initial guessed vertex parameters ($\textbf{v}_{o},\textbf{t}_{o},k_{o}$):
$h(\textbf{v}_{o}+\delta\textbf{v},\textbf{t}_{i,o}+\delta\textbf{t}_{i},k_{i,o}+\delta k_{i})_{j}\simeq h(\textbf{v}_{o},\textbf{t}_{i,o},k_{i,o})_{j}+D_{i}\,\delta\textbf{v}+E_{i}\,\delta\textbf{t}_{i}\,+F_{i}\,\delta k_{i},$ (7)
@@ III.2 Least-squares algorithm
$\chi^{2}=\sum_{i}\left(\delta q_{i}-D_{i}\delta\textbf{v}-E_{i}\delta\textbf{t}_{i}-F_{i}\delta k_{i}\right)^{T}W_{i}\left(\delta q_{i}-D_{i}\delta\textbf{v}-E_{i}\delta\textbf{t}_{i}-F_{i}\delta k_{i}\right).$ (12)
where $W=\Xi^{-1}$ is the weight matrix (see for instance BILLOIR1985115 ) and $\delta q$ is the residual at the RS:
@@
The solutions of system described in eq. 14 are:
$\delta\textbf{v}$ $=\left[\left(\sum_{i}A^{T}-\sum_{i}B^{T}_{i}C^{-1}_{i}B_{i}\right)+\left(\sum_{i}B^{T}_{i}C^{-1}_{i}G_{i}-\sum H_{i}^{T}\right)N_{1}N_{3}\right]^{-1}$ (16)
×$\times\left(\sum_{i}T_{i}-\sum_{i}B_{i}^{T}C^{-1}_{i}U_{i}+\left(\sum_{i}B^{T}_{i}C^{-1}_{i}G_{i}-\sum H_{i}^{T}\right)N_{1}N_{2}\right),$
@@
$N_{1}$ $=\left(L_{i}-G^{T}_{i}C^{-1}_{i}G_{i}\right)^{-1}\,,$ (17)
@@
$\text{Cov}(\textbf{t}_{i})$ $=C_{i}^{-1}+\left(C_{i}^{-1}B_{i}\right)\text{Cov}(\textbf{v})\left(C_{i}^{-1}B_{i}\right)^{T}+\left(C_{i}^{-1}G_{i}\right)\text{Cov}(k)\left(C_{i}^{-1}G_{i}\right)^{T}.$
III.3 Algorithm testing
Geant4 based Monte-Carlo (MC) simulations of Mu3e decays have been carried out for testing the vertex fitting described in section III.2. The procedure followed by the test was:

### v2

@@ Abstract
This paper provides a detailed description of a vertex fitting algorithm designed for precision measurements in low-energy particle physics experiments. An accurate reconstruction of low-momentum trajectories is facilitated by reducing the material budget of the detector to a few per mill of the radiation length. This decreases the multiple scattering experienced by particles within the detector, thereby enhancing vertex fitting accuracy. However, in the case of light detection systems, the intrinsic spatial resolution of the sensors imposes further constraints on the final vertex resolution that require careful consideration. The algorithm developed in this study addresses both multiple scattering and spatial resolution aspects in the context of vertex fitting, specifically tailored for light pixel detectors. Furthermore, this works presents a detailed examination of the vertex reconstruction within the low-material budget pixel detector of the Mu3e experiment.
keywords:
Vertex fitting , pixel detectors , Mu3e , Lepton Flavour Violation
[…]
organization=Department of Physics, University of Liverpool, addressline=The Oliver Lodge Laboratory, city=Liverpool, postcode=L69 7ZE, state=, country=United Kingdom
1 Introduction
@@
With the increase of the instantaneous and integrated beam luminosities, the requirements of particle physics experiments for precise tracking and vertexing detectors, with high radiation tolerance, have become more stringent e.g., [1, 2, 3, 4]. In this regard, silicon pixel sensors can provide high granularity, low material budget structures and the radiation-hardness that most experiments need, e.g., [4, 5, 6, 7]. It is important, however, that precise detection systems are developed in conjunction with equally performing analysis methods for the reconstruction of particle trajectories and decay vertices, e.g. [8, 9, 10, 11]. To this aim, several fitting algorithms have been designed and optimized over the years to deal with hit selection, pattern recognition, errors calculations and high track multiplicity (see for instance [11, 12] and references therein). The practical implementation of these methods must be tailored around the actual detector and magnetic field configuration of the experiments. This makes tracking and vertexing a topic which is in continuous evolution adapting itself to the new challenges set by upcoming experiments.
@@
This study addresses the problem of vertex fitting in the low-material budget pixel detector of the Mu3e experiment [7]. As explained in section 2, the detector design has been optimized to minimize the effects of Multiple Coulomb Scattering (MS) on particle trajectories and signal kinematics. However, in the case of light detectors like Mu3e, the intrinsic pixel resolution emerges as an additional constraint in the vertex reconstruction demanding careful consideration. The present vertex fitting takes into account both MS and pixel resolution errors, alongside any uncertainties derived from energy losses within the detector. Section 3 provides a detailed description of the algorithm whilst in section 4, a comparative study is conducted among different inclusive scenarios that encompass: (A) MS only; (B) MS and pixel resolution together; (C) all sources of errors (MS, pixel resolution and energy losses) are included.
2 The Mu3e low-material budget pixel detector
Figure 1: Scheme of the central station of the Mu3e pixel detector, side and x-y views [7].
@@ 2 The Mu3e low-material budget pixel detector
at final Branching Ratios (BR) $>10^{-16}$ [7]. The first phase of the experiment is designed to use the $\pi$E5 low-energy muon source ($10^{8}\,$z) at the Paul Scherrer Institute, where Mu3e is under commissioning. The aimed single event sensitivity of the experiment is $4$ orders of magnitude higher than previous experimental upper limits (BR $=1.0\cdot 10^{-12}$) [13] and $38$ orders of magnitude lower than theoretical Standard Model (SM) calculations (BR $=10^{-54}$), e.g., [14, 15]. However, new theoretical studies predict the existence of additional degrees of freedom beyond the SM. These extensions may lead to CLF violation potentially detectable by upcoming experiments such as Mu3e, e.g., [16, 17]. Consequently, the observation of $\mu^{+}\rightarrow e^{+}e^{-}e^{+}$ at sensitivity levels achievable in the Mu3e experiment will be a clear signature of new physics beyond the SM [17, 18].
The process in (1) yields a relatively simple decay topology with the 3 final state leptons produced at the same vertex of interaction and momentum vectors, $\vec{p}$, determined by the energy and momentum conservation for decaying muons at rest. The main background processes in Mu3e measurements are the muon internal conversion $\mu^{+}\rightarrow e^{+}e^{-}e^{+}+\nu_{e}+\bar{\nu}_{\mu}$ (BR $\approx 10^{-5}$) and the combination of one electron and 2 positrons from independent sources, e.g., one Bhabha electron plus two Michel positrons $\mu^{+}\rightarrow 2\times(e^{+}+\nu_{e}+\bar{\nu}_{\mu})+e^{-}$ [13]. The aimed single event sensitivity of $2\cdot 10^{-15}$, during phase I of the experiment, can be achieved with an energy-momentum resolution of $\lesssim 1$ MeV and by using precise vertexing and timing systems [7].
@@
Figure 1 shows the schematic of the foreseen Mu3e tracker central station which is important for vertex fitting and track reconstruction [7]. Two re-curl stations, one up-stream and one down-stream, will also be part of the final detector design. These increase the angular acceptance of the experiment and allow to measure long-armed trajectories to achieve improved momentum resolution.
The layers have cylindrical symmetry and are concentrically placed around the target, a hollow double cone made of Mylar 100 mm in length and with a base radius of 19 mm. The target is placed in a solenoid magnetic field of 1 T with the base at a minimum distance of $\approx 4$ mm from the innermost layer of the pixel tracker. Particle trajectories bend inwards following helical trajectories around the field lines possibly making multiple re-curls. Each layer of the pixel detector is sectioned in sub-elements called ladders. A ladder is a series of chips mounted on the same flex-tape. There are 8, 10, 24 and 28 ladders for layer 1, 2, 3 and 4, respectively. For instance, the innermost layer of the tracker, crucial for vertex fitting, is made of 8 ladders each one tilted by a 45 angle with respect to the neighbours. This configuration forms a 8-sided surface which extends for $\sim 12$ cm or 6 chips length, see figure 1.
@@
3 Vertex fitting in the Mu3e pixel detector
Vertex reconstruction can be accomplished in two steps: vertex finding and vertex fitting, see e.g., [11]. The former consists in grouping trajectories that have been most likely produced in the same decay process. The latter involves finding the most likely vertex coordinates $(x,y,z)_{\text{v}}$ and the initial momentum vectors of all clustered tracks. In Mu3e, vertex finding is accomplished by considering all possible combinations of two positive and one negative tracks in the detector within time frames of 64 ns. For the vertex fitting, a least-squares optimization algorithm has been developed based on the method illustrated in [9].
3.1 Track parameters and uncertainties
Particle trajectories are defined by 6 parameters ($[x,y],z,\phi,\lambda,k)$. These are the coordinates of one point along the track, the angles $\phi$ and $\lambda$ defining the direction of the tangent vector to the trajectory and the factor $k=(p/q)^{-1}$ where $q$ is the charge of the particle and $p$ is the magnitude of the momentum vector. The following relationships among the momentum components in the global Cartesian frame and the angles $\phi$ and $\lambda$ hold true:
@@ 3.1 Track parameters and uncertainties
In the present study, track parameters at a given Reference Surface (RS) along with their covariance matrix ($\Xi$): ($[x,y],z,\phi,\lambda,k,\Xi)_{\text{meas}}$ are considered as input measurements for the fit. The RS is the innermost layer of the pixel detector which is the closest one to the expected real vertex position. The coordinates $x,y$ are not independent and can be given in a single expression by using a local reference frame with center in the middle of the pixel area $(\bar{x},\bar{y},\bar{z})$ and base vectors $(\hat{u},\hat{z^{\prime}})$. The vector $\hat{z^{\prime}}$ is parallel to the global coordinate $\hat{z}$ while $\hat{u}$ is perpendicular to it and parallel to the pixel surface, see figure 2. The equations that link the local coordinates to the global ones are:
@@
where the angle $\gamma$ is the orientation of the pixel with respect to the $x$ axis of the global reference frame. Following the equations in 3, the track parameters at the RS become $(u,z,\phi,\lambda,k)$  [^1] The coordinate $z^{\prime}$ can be replaced by $z$ given that the constant $\bar{z}$ in eq. 3 does not contribute to the calculations of the derivatives and residuals carried out in section 3.2.
Figure 2: Sketch of the RS considered in this study. The axes of the global (local) reference frame are drawn in black (green); the $\hat{z^{\prime}}$ axis perpendicular to the $x-y$ plane is not shown. The parameters (${x,y},z)_{\text{meas}}$ are placed in the geometrical centre of the pixel surface $\square$. The fit parameters ($u,z,\phi,\lambda,k)_{\text{fit}}$ are obtained by forward propagating vertex parameters ($x,y,z,\phi,\lambda,k)_{\text{v}}$ via the function $h=h(\textbf{v}=(x,y,z)_{\text{v}},\textbf{t}(\phi_{\text{v}},\lambda_{\text{v}}),k_{\text{v}}))$.
The covariance matrix can be written as the sum of two terms, i.e., $\Xi_{\text{meas}}=\Xi_{\text{track}}+\Xi_{\text{RS}}$. The term $\Xi_{\text{track}}$ accounts for the uncertainties accumulated during track fitting while $\Xi_{\text{RS}}$ accounts for MS and pixel resolution at the RS. Pixel resolution contributes to the smearing of the hit position by $\sigma_{l}=l/\sqrt{12}$, where $l$ is the length of the pixel side. The MS changes the direction of the track upon crossing the RS. The tilt can be approximated by a Gaussian distribution with zero mean and standard deviation $\theta_{\text{MS}}$ [20]:
$\theta_{\text{MS}}=13.6\,\left[\text{MeV/c}\right]\,\frac{q}{p}\sqrt{\frac{d}{\beta^{2}\text{X}_{o}}}\left(1+0.038\,\text{ln}\left(\frac{d}{\text{X}_{o}}\right)\right)\,.$ (4)
In the previous equation $\beta$ is the relativistic velocity and $d$ is the distance travelled by the particle in the RS. From eq. 4, the changes in $\phi$ and $\lambda$ due to MS, $\sigma_{\phi}$ and $\sigma_{\lambda}$, can be obtained by projecting $\theta_{\text{MS}}$ onto the transverse and longitudinal planes, respectively, see e.g., [21]:
@@
In addition to MS and pixel resolution, an error $\sigma_{k}$ on the track curvature (or $k$) is introduced in $\Xi_{\text{meas}}$ to account for the energy lost by electrons and positrons in the detector, see e.g., [12]. Although energy losses in the Mu3e RS are negligible, they are nevertheless included in the present study to provide a complete treatment of the errors. Without loss of generality, the covariance matrix in this study is written as $\Xi_{\text{meas}}=\Xi_{\text{pixel}}+\Xi_{\text{MS}}+\Xi_{\text{$\Delta$E}}=$ diag$\left\{\sigma_{u}^{2},\sigma_{z}^{2},\sigma_{\phi}^{2},\sigma_{\lambda}^{2},\sigma_{k}^{2}\right\}$ with $\sigma_{u}=\sigma_{z}=\sigma_{l}$.
3.2 Least-squares algorithm
In the vertex fitting, a map that links the track vertex parameters ($x,y,z,\phi^{i},\lambda^{i},k^{i})_{\text{v}}$ to the parameters at the RS ($u^{i},z^{i},\phi^{i},\lambda^{i},k^{i})_{\text{fit}}$ is needed, as shown in figure 2. In what follows, $i=1,2,3$ for the ($e^{-},e^{+},e^{+}$) in a Mu3e decay. Starting from the analytical expression of a helical trajectory of a charged particle in a magnetic field, this map can be written as: $h_{j}=h(\textbf{v},\textbf{t},k_{\text{v}})_{j}$, where $\textbf{v}=(x,y,z)_{\text{v}}$, $\textbf{t}=(\phi,\lambda)_{\text{v}}$ and $h_{j}=(u^{i},z^{i},\phi^{i},\lambda^{i},k^{i})_{\text{fit}}$ for $j=1,2,...,5$, see A for the analytic expressions of $h_{j}$. In first approximation, this function can be linearised near some initial guessed vertex parameters ($\textbf{v}_{o},\textbf{t}_{o},k_{o}$):
$h(\textbf{v}_{o}+\delta\textbf{v},\textbf{t}_{i,o}+\delta\textbf{t}_{i},k_{i,o}+\delta k_{i})_{j}\simeq$ (7)
$h(\textbf{v}_{o},\textbf{t}_{i,o},k_{i,o})_{j}+D_{i}\,\delta\textbf{v}+E_{i}\,\delta\textbf{t}_{i}\,+F_{i}\,\delta k_{i},$
@@ 3.2 Least-squares algorithm
$\chi^{2}$ $=\sum_{i}\Delta^{T}W_{i}\Delta\,,$ (12)
$\Delta$ $=\delta q_{i}-D_{i}\delta\textbf{v}-E_{i}\delta\textbf{t}_{i}-F_{i}\delta k_{i}\,.$
where $W=\Xi^{-1}$ is the weight matrix (see for instance [22]) and $\delta q$ is the residual at the RS:
@@
The solutions of the system described in eq. 14 are:
$\delta\textbf{v}$ $=\left[\left(\sum_{i}A^{T}-\sum_{i}B^{T}_{i}C^{-1}_{i}B_{i}\right)-MN_{1}N_{3}\right]^{-1}$ (16)
×$\times\left(\sum_{i}T_{i}-\sum_{i}B_{i}^{T}C^{-1}_{i}U_{i}-MN_{1}N_{2}\right),$
@@
$M$ $=\left(\sum H_{i}^{T}-\sum_{i}B^{T}_{i}C^{-1}_{i}G_{i}\right)\,,$ (17)
$N_{1}$ $=\left(L_{i}-G^{T}_{i}C^{-1}_{i}G_{i}\right)^{-1}\,,$
@@
$\text{Cov}(\textbf{t}_{i})=$ $C_{i}^{-1}+\left(C_{i}^{-1}B_{i}\right)\text{Cov}(\textbf{v})\left(C_{i}^{-1}B_{i}\right)^{T}+$
$\left(C_{i}^{-1}G_{i}\right)\text{Cov}(k)\left(C_{i}^{-1}G_{i}\right)^{T}.$
3.3 Algorithm testing
Geant4 based Monte-Carlo (MC) simulations of Mu3e decays have been carried out for testing the vertex fitting described in section 3.2. The procedure followed by the test was: